# Limit Computation Over Posets via Minimal Initial Functors

### v1

@@ 5.3. Size Bounds
$D_{1}=\{(i,k-i,a,b)\mid(a,b)<(j,k-j)\},$
$D_{2}=\{(a,b,j,k-j)\mid(a,b)<(i,k-i)\},$
we have $\operatorname{dn}(t_{ij},U)=D_{1}\sqcup D_{2}$. Moreover, every point in $D_{1}$ is incomparable to every point in $D_{2}$. Any path in $\operatorname{dn}(t_{ij},U)$ connecting a point in $D_{1}$ to a point in $D_{2}$ must contain a pair of comparable points $p,p^{\prime}$, with $p\in D_{1}$ and $p^{\prime}\in D_{2}$. Thus, no such path can exist, which implies that $t_{ij}$ is essential. Therefore, $U$ has at least $(k+1)^{2}$ essential elements, so by Theorems˜5.5 and 5.13, we have $|I_{U}|=\Theta(k^{2})$, as claimed.
@@
Suppose $P\subseteq Q$ is an initial scaffold of an interval $Q\subseteq\mathbb{N}^{d}$. Note that $\operatorname{dn}(p,Q)$ as at most $d$ components for each $p\in Q$. Thus, for each $p\in P$, there are at most $d$ relations in $P$ of the form $m<p$. Hence, to prove the result, it suffices to show that $|\operatorname{Ob}P|=|I_{Q}|$ satisfies the bounds of the theorem. By Lemma˜5.8, we have $|I_{Q}|\leq|I_{\operatorname{up}[Q]}|$. Since $\operatorname{up}[Q]\subseteq\mathbb{N}^{d}$ is an upset with the same minima as $Q$, Corollary˜5.7 and Lemma˜5.13 together imply that
@@ 6.1. Algorithm for Finite Posets
We next give an algorithm to compute the initial scaffold of a finite poset, thereby proving Theorem˜3.13. Recall the statement:
@@
If we choose an initial scaffold for each component of $Q$, then the union of these initial scaffolds is an initial scaffold for $Q$. The components of $Q$ can be computed in time $O(|V|+|E|)=O(|V||E|)$ by depth-first search on the Hasse diagram. It therefore suffices to treat the case of a connected poset $Q$. Assuming $Q$ is connected, the algorithm to compute an initial scaffold $P\subseteq Q$ then proceeds as follows: We refine the partial order on $Q$ to a total order via the standard topological sorting algorithm in time $O(|V|+|E|)=O(|E|)$. We process the elements of $Q$ in increasing order. When we visit $q\in Q$, we compute the components of $\operatorname{dn}(q,Q)$ using depth-first search in time $O(|E|)$. If the number of components is not 1, then we add $q$ to $P$ and, for each component $A$ of $\operatorname{dn}(q,Q)$, we add a relation $m<q$ in $P$ for some choice of minimum $m\in A$. These minima can be identified while computing the components of $\operatorname{dn}(q,Q)$ by counting in-degrees of vertices. It is clear that the total time cost of the algorithm is $O(|V||E|)$. ∎
@@ 6.2. Algorithm for Intervals in $\mathbb{N}^{3}$
The rest of this section is devoted to the proof of Theorem˜3.17. Let us recall the statement:
@@
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=s$, we can compute an initial scaffold of $Q$ in time
@@
    $O(s^{4})$ for $d>3$.
In the case that $Q$ is finite, these bounds also hold when $Q$ is instead specified by its extrema, with $s$ the total number of these.
We first prove Theorem˜3.17 (i) by giving an algorithm that computes an initial scaffold of an interval $Q\subseteq\mathbb{N}^{3}$ and bounding its complexity. The algorithm and its analysis specialize to intervals $Q\subseteq\mathbb{N}^{2}$ by embedding $\mathbb{N}^{2}$ into $\mathbb{N}^{3}$ via the map $(x,y)\mapsto(x,y,0)$.
Letting $U=\operatorname{up}[Q,\mathbb{N}^{3}]$, our algorithm first computes an initial scaffold $P$ of $U$ and then computes $P\cap Q$. According to the following result, this is an initial scaffold of $Q$.
@@
The computation of $P$ proceeds by iterating through 2-D coordinate slices of $U$. To elaborate, for any upset $I\subseteq\mathbb{N}^{3}$ and $z\in\mathbb{N}$, the coordinate slice
@@
is also an upset. To compute $P$, we iterate through values of $z\in\mathbb{N}$ in increasing order, maintaining the sets of minima $M_{U_{z}}$ sorted by $x$-coordinate. When we pass from level z1z-1 to level $z$, we update $M_{U_{z-1}}$ to obtain $M_{U_{z}}$ and record all elements of $E_{U}$ of the form $e=(x,y,z)$. For each such $e$, we compute the relations $m<e$ to be added to $P$.
The computation of $P\cap Q$ from $P$ is done in a similar but much simpler way, also by iterating through 2-D coordinate slices. As we explain below, whether the poset $Q$ is specified by its upset presentation or by its extrema, we use essentially the same algorithm to compute $P\cap Q$ from $P$.
We next explain in detail how we compute $P$. We begin by introducing some notation.
@@
  • For $m=(m_{1},m_{2})\in\mathbb{N}^{2}$ and $y\in\mathbb{N}$, we identify $(m,y)$ with $(m_{1},m_{2},y)\in\mathbb{N}^{3}$.
  • For $m\in M_{U_{z}}$, let $m^{\downarrow}=(m,y)\in\mathbb{N}^{3}$, where $y\in\mathbb{N}$ is the unique value such that $(m,y)\in M_{U}$.
  • Write $M_{U_{z}}$ in increasing order of $x$-coordinate as
@@
  • Let $W^{z}$ denote the set of elements of the form $(m_{i}\vee m_{i+1},z)$ such that for some $j\in\{i,i+1\}$, both of the following conditions hold:
@@
Our approach to computing an initial scaffold $P$ of $U$ rests on the following proposition:
@@
      If $p=(m_{i}\vee m_{i+1},z)\in W^{z}$, then $P$ contains the relations $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$. (Note that either $m_{i}^{\downarrow}=(m_{i},z)$ or $m_{i+1}^{\downarrow}=(m_{i+1},z)$).
@@
      If $p=(m,z)\in X^{z}$, then $P$ contains the relation $m^{\downarrow}<p$. If in addition $p\not\in W^{z}$, then $P$ also contains exactly one relation $(a,z)<p$, where $(a,z)$ is an arbitrary element of $M_{U}$ with $a<m$.
@@
$U=\{(0,6,0),(1,5,0),(3,4,0),(4,2,0),(5,0,0),(1,3,1),(2,2,1),(4,1,1)\}.$
Then as illustrated in Fig.˜3, we have
@@
Note that $(4,2,1)\in W^{1}\cap X^{1}$. Also note that $(1,6,1)\not\in W^{1}$, because the second condition in the definition of $W^{z}$ is not satisfied.
@@
In what follows, all downsets are taken in $U$. Consider an element $p=(m_{i}\vee m_{i+1},z)\in W^{z}$, and let $j\in\{i,i+1\}$ be the index satisfying the condition in the definition of $W^{z}$ for $p$. Assume without loss of generality that $j=i$, so that $(m_{i},z)\in M_{U}$. To see that $p\in E_{U}$, it suffices to show that $(m_{i},z)$ and $(m_{i+1},z)$ belong to different components of $\operatorname{dn}(p)$. Clearly, $(m_{i},z),(m_{i+1},z)\in\operatorname{dn}(p)$. The component of $\operatorname{dn}(p)$ containing $m_{i}$ is
$\{(k,(m_{i})_{2},z)\mid(m_{i})_{1}\leq k<(m_{i+1})_{1}\}.$
@@
If $p=(m,z)\in X^{z}$, then $m\in M_{U_{z-1}}$. Moreover, there is some $a\in M_{U_{z}}$ with $a<m$. Note that $(m,z-1),(a,z)\in\operatorname{dn}(p)$. Since $m\in M_{U_{z-1}}$, it must be that $U$ contains no points of the form (b,z1)(b,z-1) with $b<m$. Thus, (m,z1)(m,z-1) and $(a,z)$ lie in different components of $\operatorname{dn}(p)$. Hence, $p$ is essential.
Conversely, suppose that $p=(p_{1},p_{2},z)$ is essential. Then since $U$ is an interval, at least two of the three vectors $\{p-\mathbf{e}_{1},p-\mathbf{e}_{2},p-\mathbf{e}_{3}\}$ lie in distinct components of $\operatorname{dn}(p)$. We first consider the case where $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ lie in distinct components of $\operatorname{dn}(p)$, which we denote as $C_{1}$ and $C_{2}$, respectively. We will show that in this case, $p\in W^{z}$. Since $C_{1}$ and $C_{2}$ are disjoint, we have $p-\mathbf{e}_{1}-\mathbf{e}_{2}\not\in\operatorname{dn}(p)$. Therefore,
$C_{1}$ $\subseteq\{(a,p_{2},c)\mid a\leq p_{1},\ c\leq z\},$
$C_{2}$ $\subseteq\{(p_{1},b,c)\mid b\leq p_{2},\ c\leq z\}.$
@@
$m_{i}$ $=\min\{(a,p_{2})\mid(a,p_{2},z)\in C_{1}\},$
$m_{i+1}$ $=\min\{(p_{1},b)\mid(p_{1},b,z)\in C_{2}\}.$
Moreover, since $C_{1}$ and $C_{2}$ are disjoint, they cannot both contain $p-\mathbf{e}_{3}$. Assume without loss of generality that $p-\mathbf{e}_{3}\not\in C_{1}$. We will show that $p$ satisfies the definition of an element of $W^{z}$, for $j=i$. Since $U$ is an interval, we have
$C_{1}\subseteq\{(a,p_{2},z)\mid a<p_{1}\}.$
@@
Next, suppose that $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ do not lie in distinct components of $\operatorname{dn}(p)$, i.e., either $\{p-\mathbf{e}_{1},p-\mathbf{e}_{2}\}\not\subseteq\operatorname{dn}(p)$ or else $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ lie in the same component of $\operatorname{dn}(p)$ (in which case $p-\mathbf{e}_{1}-\mathbf{e}_{2}\in\operatorname{dn}(p)$). Assume without loss of generality that $p-\mathbf{e}_{1}\in\operatorname{dn}(p)$. Then $p-\mathbf{e}_{3}$ lies in a separate component of $\operatorname{dn}(p)$; call the respective components $U_{1}$ and $U_{3}$. Note that neither $p-\mathbf{e}_{3}-\mathbf{e}_{1}\in\operatorname{dn}(p)$ nor $p-\mathbf{e}_{3}-\mathbf{e}_{2}\in\operatorname{dn}(p)$, since otherwise we would have $U_{1}=U_{3}$, a contradiction. It follows that $p-\mathbf{e}_{3}\in M_{U_{z-1}}$, and since $p-\mathbf{e}_{1}\in U$, we also have that $p\not\in M_{U_{z}}$. Hence $p\in X^{z}$. This completes the proof of (i).
@@
To prove (ii), suppose $p=(m_{i}\vee m_{i+1},z)\in W^{z}$. The proof of (i) shows that $(m_{i},z)$ and $(m_{i+1},z)$ lie in separate components of $\operatorname{dn}(p)$, so $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$ are valid choices of relations in $P$. If in addition, $p\not\in X^{z}$, then either $p-\mathbf{e}_{3}\not\in\operatorname{dn}(p)$ or $p-\mathbf{e}_{3}-e_{i}\in\operatorname{dn}(p)$ for some $j\in\{0,1\}$. In either case, $\operatorname{dn}(p)$ has two components, so no other relations with upper element $p$ need to be added to $P$.
@@
Suppose $p=(m,z)\in X^{z}$, then since $p-\mathbf{e}_{3}\in\operatorname{dn}(p)$ and there is no $b<m$ with $(b,z-1)\in\operatorname{dn}(p)$, the interval property of $U$ implies that there is a component of $\operatorname{dn}(p)$ lying in the set $\{(m,y)\in\mathbb{N}^{3}\mid y<z$. It follows that $P$ must contain the relation $m^{\downarrow}<p$. If in addition $p\not\in W^{z}$, then $\operatorname{dn}(p)$ has exactly one component not containing $p-\mathbf{e}_{3}$, and all elements of this component have third coordinate $z$. Therefore, for $(a,z)$ an arbitrary element of $M_{U}$ with $a<m$, the relation $(a,z)<p$, is also valid in $P$, and since $\operatorname{dn}(p)$ has two components, no other relations with upper element $p$ need to be added to $P$.

### v2

@@ 5.3. Size Bounds
$D_{1}=\{(i,k-i,a,b)\in\mathbb{N}^{4}\mid(a,b)<(j,k-j)\},$
$D_{2}=\{(a,b,j,k-j)\in\mathbb{N}^{4}\mid(a,b)<(i,k-i)\},$
every point in $D_{1}$ is incomparable to every point in $D_{2}$. Thus, $\operatorname{dn}(e_{ij},U)=D_{1}\sqcup D_{2}$, which implies that $e_{ij}$ is essential. Therefore, $|E_{U}|\geq(k+1)^{2}$, so by Theorems˜5.5 and 5.12, we have $|E_{U}|=\Theta(k^{2})$, as claimed.
@@
Suppose $P\subseteq Q$ is an initial scaffold of an interval $Q\subseteq\mathbb{N}^{d}$. Note that $\operatorname{dn}(p,Q)$ has at most $d$ components for each $p\in Q$. Thus, for each $p\in P$, there are at most $d$ relations in $P$ of the form $m<p$. Hence, $|P|=O(|\operatorname{Ob}P|)$. Therefore, to prove the result, it suffices to show that $|\operatorname{Ob}P|=|I_{Q}|$ satisfies the bounds of the theorem. By Lemma˜5.8, we have $|I_{Q}|\leq|I_{\operatorname{up}[Q]}|$. Since $\operatorname{up}[Q]\subseteq\mathbb{N}^{d}$ is an upset with the same minima as $Q$, Corollary˜5.7 and Lemma˜5.13 together imply that
@@ 6.1. Algorithm for Finite Posets
We next prove Theorem˜3.13 by giving an algorithm to compute the initial scaffold of a finite poset. Recall the theorem statement:
@@
The components of $Q$ can be computed in time $O(|V|+|E|)=O(|V||E|)$ by depth-first search on the Hasse diagram. Given an initial scaffold for each component of $Q$, the union of these is an initial scaffold of $Q$. It therefore suffices to treat the case of a connected poset $Q$.
Assuming $Q$ is connected, our algorithm to compute an initial scaffold $P\subseteq Q$ proceeds as follows: We refine the partial order on $Q$ to a total order via the standard topological sorting algorithm in time $O(|V|+|E|)=O(|E|)$. We process the elements of $Q$ in increasing order. When we visit $q\in Q$, we use depth-first search to compute the components of $\operatorname{dn}(q,Q)$, as well as a choice of minimum $m_{A}\in A$ for each component $A$ of $\operatorname{dn}(q,Q)$; this takes time $O(|E|)$. If the number of components is not 1, then we add $q$ to $P$ and we also add the relation $m_{A}<q$ to $P$ for each component $A$ of $\operatorname{dn}(q,Q)$. It is clear that the total time cost of this algorithm is $O(|V||E|)$. ∎
@@ 6.2. Algorithm for Intervals in $\mathbb{N}^{3}$
The rest of this section is devoted to the proof of Theorem˜3.16. Let us recall the statement:
@@
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=n$, we can compute an initial scaffold of $Q$ in time
@@
    $O(n^{4})$ for $d>3$.
In the case that $Q$ is finite, these bounds also hold when $Q$ is instead specified by its extrema, with $n$ the total number of these.
Our approach is based on the following observation:
@@
We prove Theorem˜3.16 (i) by giving an algorithm that computes an initial scaffold of an interval $Q\subseteq\mathbb{N}^{3}$ in time $O(n\log n)$. The algorithm and its analysis specialize to intervals $Q\subseteq\mathbb{N}^{2}$ by embedding $\mathbb{N}^{2}$ into $\mathbb{N}^{3}$ via the map $(x,y)\mapsto(x,y,0)$.
Letting $U=\operatorname{up}[Q,\mathbb{N}^{3}]$, our algorithm first computes an initial scaffold $P$ of $U$ and then computes $P\cap Q$. The computation of $P$ proceeds by iterating through planar slices of $U$. To elaborate, for any upset $I\subseteq\mathbb{N}^{3}$ and $z\in\mathbb{N}$, the planar slice
@@
is also an upset. To compute $P$, we iterate through values of $z\in\mathbb{N}$ in increasing order, maintaining the sets of minima $M_{U_{z}}$. When we pass from level z1z-1 to level $z$, we update $M_{U_{z-1}}$ to obtain $M_{U_{z}}$ and record all points of $E_{U}$ of the form $e=(x,y,z)$. For each such $e$, we also compute the relations $m<e$ to be added to $P$. To carry out these computations, we leverage a concrete levelwise description of $P$, given in Proposition˜6.3 below.
The computation of $P\cap Q$ from $P$ is done in a similar but much simpler way, also by iterating through planar slices. As we explain below, whether the poset $Q$ is specified by its upset presentation or by its extrema, we use essentially the same approach to compute $P\cap Q$ from $P$.
@@
  • (i)
    For $m=(m_{1},m_{2})\in\mathbb{N}^{2}$ and $t\in\mathbb{N}$, we identify $(m,t)$ with $(m_{1},m_{2},t)\in\mathbb{N}^{3}$.
  • (ii)
    For $m\in M_{U_{z}}$, let $m^{\downarrow}=(m,t)$, where $t\in\mathbb{N}$ is the unique value such that $(m,t)\in M_{U}$.
  • (iii)
    We write $M_{U_{z}}$ in increasing order of $x$-coordinate as
@@
  • (iv)
    Let $W^{z}$ denote the set of points of the form $(m_{i}\vee m_{i+1},z)$ such that for some $j\in\{i,i+1\}$, both of the following conditions hold:
@@
      If $p=(m_{i}\vee m_{i+1},z)\in W^{z}$, then $P$ contains the relations $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$.
@@
      If $p=(m,z)\in X^{z}$, then $P$ contains the relation $m^{\downarrow}<p$. If in addition $p\not\in W^{z}$, then $P$ also contains exactly one relation $(l,z)<p$, where $(l,z)$ is an arbitrary point of $M_{U}$ with $l<m$.
@@
$U=\langle(0,6,0),(1,5,0),(3,4,0),(4,2,0),(5,0,0),(1,3,1),(2,2,1),(4,1,1)\rangle.$
Then, as illustrated in Fig.˜3, we have
@@
Note that $W^{1}\cap X^{1}=\{(4,2,1)\}$. Also note that $(1,6,1)\not\in W^{1}$, because the second condition in the definition of $W^{z}$ is not satisfied.
@@
In what follows, all downsets are taken in $U$. Consider a point $p=(m_{i}\vee m_{i+1},z)\in W^{z}$. Recalling the definition of $j\in\{i,i+1\}$ in ˜6.2 (iv), assume without loss of generality that $j=i$, so that $(m_{i},z)\in M_{U}$. To see that $p\in E_{U}$, it suffices to show that $(m_{i},z)$ and $(m_{i+1},z)$ belong to different components of $\operatorname{dn}(p)$. Clearly, $(m_{i},z),(m_{i+1},z)\in\operatorname{dn}(p)$. Note that $\operatorname{dn}(p)$ contains no point of the form $(a,b,z)$ with $a<(m_{i+1})_{1}$ and $b<(m_{i})_{2}$, since otherwise $m_{i}$ and $m_{i+1}$ would not be consecutively ordered by $x$-coordinate in $M_{U_{z}}$. Moreover, by the definition of $W^{z}$, there is no $x\in M_{U_{z-1}}$ with $m_{i}<x<m_{i}\vee m_{i+1}$, which implies that $\operatorname{dn}(p)$ contains no point of the form $(a,(m_{i})_{2},c)$ with $a<(m_{i+1})_{1}$ and $c<z$. Therefore, the component of $\operatorname{dn}(p)$ containing $m_{i}$ is
$\{(a,(m_{i})_{2},z)\mid(m_{i})_{1}\leq a<(m_{i+1})_{1}\}.$
@@
If $p=(m,z)\in X^{z}$, then $m\in M_{U_{z-1}}$ and moreover, there is some $l\in M_{U_{z}}$ with $l<m$. Note that $(m,z-1),(l,z)\in\operatorname{dn}(p)$. Since $m\in M_{U_{z-1}}$, it must be that $U$ contains no points of the form $(l^{\prime},z-1)$ with $l^{\prime}<m$. Thus, (m,z1)(m,z-1) and $(l,z)$ lie in different components of $\operatorname{dn}(p)$. Hence, $p$ is essential.
Conversely, suppose that $p=(p_{1},p_{2},z)$ is essential. Since $U$ is an upset, every component of $\operatorname{dn}(p)$ contains a point of the form $p-\mathbf{e}_{t}$ for some $t$. Thus, since $p$ is essential, at least two of the three vectors $\{p-\mathbf{e}_{1},p-\mathbf{e}_{2},p-\mathbf{e}_{3}\}$ lie in distinct components of $\operatorname{dn}(p)$. We first consider the case where $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ lie in distinct components of $\operatorname{dn}(p)$, denoted $A_{1}$ and $A_{2}$, respectively. We now show that in this case, $p\in W^{z}$. Since $A_{1}$ and $A_{2}$ are disjoint, it must be that $p-\mathbf{e}_{1}-\mathbf{e}_{2}\not\in\operatorname{dn}(p)$. Therefore,
$A_{1}$ $\subseteq\{(a,p_{2},c)\mid a\leq p_{1},\ c\leq z\},$
$A_{2}$ $\subseteq\{(p_{1},b,c)\mid b\leq p_{2},\ c\leq z\}.$
@@
$m_{i}$ $=\min\{(a,p_{2})\mid(a,p_{2},z)\in A_{1}\},$
$m_{i+1}$ $=\min\{(p_{1},b)\mid(p_{1},b,z)\in A_{2}\}.$
Moreover, since $A_{1}$ and $A_{2}$ are disjoint, they cannot both contain $p-\mathbf{e}_{3}$. Assume without loss of generality that $p-\mathbf{e}_{3}\not\in A_{1}$. Since $U$ is an upset, we must have
$A_{1}\subseteq\{(a,p_{2},z)\mid a<p_{1}\}.$
@@
Next, suppose that $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ do not lie in distinct components of $\operatorname{dn}(p)$, i.e., either $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ lie in the same component of $\operatorname{dn}(p)$, in which case $p-\mathbf{e}_{1}-\mathbf{e}_{2}\in\operatorname{dn}(p)$, or else $\{p-\mathbf{e}_{1},p-\mathbf{e}_{2}\}\not\subseteq\operatorname{dn}(p)$. Assume without loss of generality that $p-\mathbf{e}_{1}\in\operatorname{dn}(p)$. Then $p-\mathbf{e}_{3}$ lies in a separate component of $\operatorname{dn}(p)$; call the respective components $A_{1}$ and $A_{3}$. Note that neither $p-\mathbf{e}_{3}-\mathbf{e}_{1}\in\operatorname{dn}(p)$ nor $p-\mathbf{e}_{3}-\mathbf{e}_{2}\in\operatorname{dn}(p)$, since otherwise we would have $A_{1}=A_{3}$, a contradiction. It follows that $p-\mathbf{e}_{3}\in M_{U_{z-1}}$, and since $p-\mathbf{e}_{1}\in U$, we also have that $p\not\in M_{U_{z}}$. Hence $p\in X^{z}$. This completes the proof of (i).
@@
To prove (ii), suppose $p=(m_{i}\vee m_{i+1},z)\in W^{z}$. The proof of (i) shows that $(m_{i},z)$ and $(m_{i+1},z)$ lie in separate components of $\operatorname{dn}(p)$, so $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$ are valid choices of relations in $P$. If in addition $p\not\in X^{z}$, then $p-\mathbf{e}_{3}\not\in\operatorname{dn}(p)$, so $\operatorname{dn}(p)$ has two components. Thus, no other relations with upper point $p$ need to be added to $P$.
@@
If $p=(m,z)\in X^{z}$, then $p-\mathbf{e}_{3}\in\operatorname{dn}(p)$ and there is no $l<m$ with $(l,z-1)\in\operatorname{dn}(p)$. Thus, one component of $\operatorname{dn}(p)$ is contained in the set $\{(m,c)\in\mathbb{N}^{3}\mid c<z\}$. It follows that $P$ must contain the relation $m^{\downarrow}<p$. If in addition $p\not\in W^{z}$, then $\operatorname{dn}(p)$ has exactly one other component, all points of which have third coordinate $z$. Therefore, for $(l,z)$ an arbitrary point of $M_{U}$ with $l<m$, the relation $(l,z)<p$ is a valid choice of relation in $P$. Since $\operatorname{dn}(p)$ has two components, no other relations with upper point $p$ need to be added to $P$.